OpenStudy (akohl):

@Abhisar At the top of a looped section of roller coaster track, the car and rider are completely upside down. What information is necessary to calculate the minimum speed of the car that will prevent a rider from falling out of it? Assume the rider is not strapped into the car. the mass of the rider and the width of the track the radius of the loop, the mass of the rider. and the mass of the car the radius of the loop the radius of the loop and the mass of the rider

OpenStudy (abhisar):

Tell me the formula of centrifugal force ?

OpenStudy (akohl):

Is it v *r

OpenStudy (abhisar):

No! It is \(\sf\huge\frac{mv^2}{r}\)

OpenStudy (abhisar):

No in order to fulfill the given condition centrifugal force should be equal to weight of the rider.

OpenStudy (abhisar):

*now

OpenStudy (akohl):

So we need both masses and the radius of the track

OpenStudy (abhisar):

No only radius of the loop

OpenStudy (abhisar):

\(\huge\sf \frac{\cancel{m}v^2}{r}=\cancel{m}g\\\huge v^2=rg\)

OpenStudy (abhisar):

getting it ?

OpenStudy (akohl):

Yea thanks

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)